APACHE POINT OBSERVATORY SDSS 2.5M OBSERVING LOG Friday May 18, 2007 (MJD 54239) ---=== OBSERVING TEAM ===--- Kaike Pan Audrey Simmons ---=== OBSERVING PLAN ===--- Science! ---=== OBSERVING SUMMARY ===--- Clouds, close radar returns, and high humidity kept us from observing most of the night and lightning closeby kept us from completeing any imager calibrations. However, at around 08:00Z, the skies did manage to clear up and we started doing science. We took two good exposures of cart 6, legacy chunk165 plate 2754 before shutting down for high humidity. No problems to report. ---=== OBSERVING LOG ===--- Afternoon: ---------- Spectro focus set for a temperature of 8 degrees. Night: ------ Did not open at sunset due to clouds, close radar returns, and high humidity. 03:42Z - We set up for image calibrations, but waited for lightning in the area to dissapate before starting. We ended up not doing calibrations because the sky began to clear up. The humidity and dewpoint difference values were barely acceptable though. We decided to give science a try anyways. 08:09Z - Opened enclosure and mounted cart 6, legacy plate 2754 - chunk165. There were too large of fiducial corrections for the altitude axis during goToField, but guide stars were still found. 08:28Z - Took precals and started a 15 min exposure - seeing about 1.2". 08:43Z - 20 min exposure. Plate not finished. 09:05Z - Closed enclosure due to high humidity. Remained closed the rest of the night. 09:49Z - endNight rejoiced ---=== IMAGING RUN SUMMARY ===--- Run Time Stripe Lambda Last Flavor Comments Start End Begin End Frame ------------------------------------------------------------------------- 6830 00:24Z 00:43Z 100 O -48.45 -43.76 45 ignore ---=== IMAGING RUN DETAILS ===--- ---=== SKIPPY RESULTS ===--- Run Frame nFrames stars muErr muRms nuErr nuRms rot az el --------------------------------------------------------------------------- ---=== LTMATCH RESULTS ===--- Run Field nFields alt az nGood rowMean rowSig colMean colSig rot ------------------------------------------------------------------------ ---=== SPECTROSCOPY DATA SUMMARY ===--- Summary Checked (y/n): yes QA Procedures Done (y/n): yes UT Exp Time flavor comment (S/N)^2 totals ========================================== b1 r1 b2 r2 ----- sequence 44954, plate -9999 ------- 00:26 44954 0.0 bias ----- sequence 44955, plate 1830 ------- 00:46 44955 10.0 flat 00:49 44956 2.0 arc ----- sequence 44957, plate 2754 ------- 08:17 44957 10.0 flat 08:19 44958 2.0 arc 08:38 44959 901.0 target 09:01 44960 1200.0 target ---=== TELESCOPE OFFSETS AND SCALE I ===--- Time Instrument Az Alt Rot Scale pos offset pos offset pos offset ------------------------------------------------------------------------------ 08:28Z 6 2754 -52.98 -0.0020 59.20 0.0055 136.59 -0.0198 1.000050 ---=== TELESCOPE OFFSETS AND SCALE II ===--- ---=== FOCUS LOG ===--- setmir piston Temp Wind Time Inst scale M1 M2 Foc Az Alt (C) MPH Dir filt fwhm ------------------------------------------------------------------------------ 00:39Z 1 1830 1.00000 0 1248 0 121 30.4 9.9 0 116 - N/A 08:28Z 6 2754 1.00005 -494 805 60 -53 59.2 8.8 2 269 - 2 ---=== WEATHER LOG ===--- Wind Time Temp F Dewp F MPH Direction Dust DIMM Sky 00:03Z 49 44 0 258 (WSW) 735 - Cloudy 00:36Z 49 45 0 222 (SW) 683 - " 01:06Z 49 43 2 341 (NNW) 702 - " 01:36Z 49 39 7 322 (NW) 697 - " 02:08Z 48 42 5 271 (W) 692 - " 02:41Z 48 44 1 269 (W) 698 - " 03:13Z 49 35 6 280 (W) 1078 - " 03:45Z 48 39 5 263 (W) 883 - " 04:17Z 46 43 1 209 (SSW) 709 - " 04:47Z 47 44 0 279 (W) 794 - " 05:17Z 47 44 0 214 (SW) 747 - " 05:47Z 46 41 3 37 (NE) 772 - " 06:17Z 47 39 6 352 (NNW) 813 - Clear 06:48Z 48 40 1 288 (WNW) 881 - " 07:20Z 46 42 0 251 (WSW) 873 - " 07:52Z 48 37 4 315 (NW) 1209 - Cloudy 08:24Z 47 37 5 267 (W) 880 - " 08:54Z 46 42 0 259 (W) 917 - " 09:26Z 46 44 1 261 (W) 805 - " ---=== TELESCOPE STATUS ===--- 07:57Z - Began ventilation. Doors and louvers opened, fans on. Opened enclosure. 09:05Z - Enclosure is on, doors and louvers are closed, and fans are off. No interlocks bypassed. Status at: 09:49Z Telescope stowed at: 30,121,00 mount Instrument mounted: Cartridge 6 Counterweights at: 250 lbs LN2 autofill systems: Connected and turned on. 180L LN2 dewar scales: Spectro 134.50 lbs, 16.82 psi Imager 215.30 lbs, 19.25 psi ---=== SOFTWARE USED ===--- IOP/SOP: v4_33_0 Watcher: v2_36_0 MCP: v6_4_0 TPM: tpm_v3_5_0 AstroDa: v15_13_1 TCC: TCC 2.7.2.1 August 6 2004 sdssProcedures: v2_27 SoS: v5_1_8 plate-mapper: v4_3_1 ---=== MIRROR NUMBERS ===--- PRIMARY: -------- Scale: 1.000000 MIGS TONIGHT NOMINAL Axial A 0.0090 -0.0050 Axial B 0.0080 0.0090 Axial C 0.0100 0.0160 Trans D 0.0090 0.0040 Lateral E 9.9695 9.9568 Lateral F 11.5443 11.5443 GALILS Commanded: 4600. -2000. 2650. -9300. -9800. -9800. Actual: 4600. -2000. 2650. -9300. -9800. -9800. SETMIR VALUES PriDesOrient: 0.00 0.00 0.00 0.00 0.00 PriOrient: 0.00 0.00 0.00 0.00 0.00 SECONDARY: ---------- Focus: 0.00 Air Temp.: 9.7 Alt.: 30.032492 MIGS TONIGHT NOMINAL Axial A -0.0310 -0.0240 Axial B -0.0310 -0.0280 Axial C -0.0450 -0.0370 Trans D -0.0100 -0.0040 GALILS Commanded: 1614024. 1594698. 1543063. -10550. -5800. Actual: 1614053. 1594719. 1543082. -10535. -5822. SETMIR VALUES SecDesOrient: 1257.00 20.00 -10.00 -100.00 176.64 SecOrient: 1256.98 20.00 -10.00 -100.43 176.56 ---=== PROBLEMS IN DETAIL ===--- No problems to report. Filename 02% 05% 10% 50% 90% 95% 98% ------------------- ------ ------ ------ ------ ------ ------ ------ sdR-b1-00044954.fit -5.4 -3.6 -2.6 2.7 7.8 9.0 11.1 sdR-b1-00044954.fit -5.4 -3.6 -2.6 2.7 7.8 9.0 11.1 sdR-b2-00044954.fit -7.6 -5.8 -4.6 1.6 7.5 9.0 11.4 sdR-b2-00044954.fit -7.6 -5.8 -4.6 1.6 7.5 9.0 11.4 sdR-r1-00044954.fit -6.3 -5.0 -3.0 2.0 7.9 10.0 12.0 sdR-r1-00044954.fit -6.3 -5.0 -3.0 2.0 7.9 10.0 12.0 sdR-r2-00044954.fit -7.0 -4.9 -3.7 2.5 8.8 10.0 13.0 sdR-r2-00044954.fit -7.0 -4.9 -3.7 2.5 8.8 10.0 13.0